# Bootstrapping Language Models with DPO Implicit Rewards

### v1

@@ 7 Conclusion
In this paper, we introduce DICE, a novel approach that leverages the implicit reward model from DPO to further align LLMs with human preferences. Our method stands out in the current landscape of LLM alignment research, as it uses the implicit reward model to iteratively refine the policy model. Empirical results demonstrate that DICE surpasses the closed-source language models such as Gemini Pro, despite utilizing a model with only 8B parameters. This efficiency in parameter usage does not come at the cost of additional human annotations or external reward models, underscoring the practicality and accessibility of our approach.
Appendix A Prompts Used by LLM-as-a-Judge
LLM-as-a-Judge Prompt Template Review the user’s question and the corresponding response using the additive 5-point scoring system described below. Points are accumulated based on the satisfaction of each criterion: Add 1 point if the response is relevant and provides some information related to the user’s inquiry, even if it is incomplete or contains some irrelevant content. Add another point if the response addresses a substantial portion of the user’s question, but does not completely resolve the query or provide a direct answer. Award a third point if the response answers the basic elements of the user’s question in a useful way, regardless of whether it seems to have been written by an AI Assistant or if it has elements typically found in blogs or search results. Grant a fourth point if the response is clearly written from an AI Assistant’s perspective, addressing the user’s question directly and comprehensively, and is well-organized and helpful, even if there is slight room for improvement in clarity, conciseness or focus. Bestow a fifth point for a response that is impeccably tailored to the user’s question by an AI Assistant, without extraneous information, reflecting expert knowledge, and demonstrating a high-quality, engaging, and insightful answer. User: {instruction} Response: {response} After examining the user’s instruction and the response, provide brief step-by-step justifications and conclude with a score between 0 to 5. You must follow the format below: Evaluation: <evaluation> Score: <score>
Figure 4: We follow the prompt template used by Yuan et al. [28] to use LLMs to judge model responses and construct paired dataset for further preference tuning.
Appendix B Optimization for Length Regularized Reward Shaping
@@
Appendix C Extended Ablation Study
@@
Figure 6: AlpacaEval 2.0 LC Win Rate across different experience replay ratio ($\gamma$) in the Llama3 setting.
@@
In the Llama3 setting, we also conduct a coarse sweeping for the experience ratio $\gamma$, and present the AlpacaEval 2.0 LC win rate in Figure 6 for two self-alignment rounds. We observe similar trends to those in the Zephyr setting, which further justify the effectiveness of the proposed experience replay: it helps to keep a balance between the more on-policy generated data and the curated offline data. The best identified value of the mixture ratio is $\gamma=0.1$.

### v2

@@ 7 Conclusion
[…]
Ethics statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Benefits of on-policy sampling
[…]
Appendix B Experiment beyond two iterations
We conduct DICE and LLM-as-a-judge baseline (the strongest baseline) for three iterations, evaluating them on AlpacaEval 2 benchmark (see Table 6).
From the table, we observe a drop in LC win rate at the third iteration for both our approach and the LLM-as-a-Judge baseline. This is a known challenge in the literature; both self-rewarding language model (Yuan et al., 2024) and SPPO (Wu et al., 2024) show successful improvement only up to three iterations, which corresponds to two iterations in our approach, as our method begins after one initial iteration of DPO training. Additionally, we use a fixed prompt set in every iteration, whereas other works employ different prompt sets across iterations. We hypothesize that using different prompt sets across iterations could help mitigate the performance drop observed in the third iteration.
Table 6: Results of AlpacaEval 2 across two base models over three iterations. LC and WR denote length-controlled and raw win rate in percentage (%) respectively.
Method zephyr-7B-beta Llama-3-8B-DPO
AlpacaEval 2 AlpacaEval 2
LC WR LC WR
Base 12.69 10.71 18.20 15.50
LLM-as-a-Judge Iter 1 15.36 17.81 20.30 21.31
LLM-as-a-Judge Iter 2 14.14 17.89 21.80 22.42
LLM-as-a-Judge Iter 3 13.95 17.23 16.86 19.78
DICE Iter 1 19.03 17.67 25.08 25.77
DICE Iter 2 20.71 20.16 27.55 30.99
DICE Iter 3 14.02 15.30 20.61 26.18
    We note that the results of Llama-3-8B-DPO base are obtained from Meng et al. (2024).
Appendix C Optimization for length regularized reward shaping
@@
Appendix D Extended ablation study
@@
Figure 5: AlpacaEval 2 LC Win Rate across different experience replay ratio ($\gamma$) in the Llama3 setting.
@@
Effects of experience replay with Llama3 backbone model. In the Llama3 setting, we also conduct a coarse sweeping for the experience ratio $\gamma$, and present the AlpacaEval 2 LC win rate in Figure 5 for two self-alignment rounds. We observe similar trends to those in the Zephyr setting, which further justify the effectiveness of the proposed experience replay: it helps to keep a balance between the more on-policy generated data and the curated offline data. The best identified value of the mixture ratio is $\gamma=0.1$.
Comparison between Length-Regularized DPO and LR reward shaping. The length-regularized DPO proposed by Park et al. (2024) incorporates a penalty term into the DPO objective to explicitly penalize length exploitation. In contrast, our approach mitigates this issue by applying reward shaping during preference dataset construction. Since both methods target the same problem, we compare their performance in the Zephyr setting with $\gamma=0$. To clarify notation, we denote the length regularization coefficient in Park et al. (2024) as $\lambda$. Length-regularized DPO uses the generated dataset $\mathcal{D}(0)$ without reward shaping, and $\lambda$ is hypertuned from $\{0.02,0.05\}$ as suggested in length-regularized DPO paper. In comparison, our method employs the dataset $\mathcal{D}(\alpha^{\star})$. All other settings remain consistent across both approaches.
As shown in Table 7, we observe that our approach achieves better performance while effectively mitigating the length exploitation. In contrast, regularized DPO exhibits a quality-length trade-off, where larger regularization coefficients successfully control response length but lead to a decline in response quality.
Table 7: Results of AlpacaEval 2 comparing techniques for mitigating length exploitation. LC and WR denote length-controlled and raw win rate in percentage (%) respectively. Len denotes the average string length of the model response.
Method $\lambda$ AlpacaEval 2
LC WR Len
No mitigation - 13.32 15.37 2570
Regularized DPO 0.02 16.30 19.16 2629
0.05 16.03 16.14 2030
Ours - 18.88 19.31 2109
PPO with Implicit Reward. To assess whether implicit rewards can be effectively utilized in reward model-based methods such as PPO (Schulman et al., 2017), we conduct experiments using PPO with an implicit reward model. Specifically, the zephyr-7B-beta model and its SFT model are used to construct the implicit reward model. Additionally, zephyr-7B-beta serves as the base model to be optimized. We adopt the implementation in OpenRLHF (Hu et al., 2024), adhering to the recommended training parameters. Evaluation results on AlpacaEval 2 are presented in Table 8.
The results indicate a slight improvement in length-controlled win rate for PPO, although the improvement is less pronounced compared to DPO. While PPO’s performance can potentially be enhanced with additional hyperparameter tuning, such efforts would require substantial resources and time. We leave this for future work.
Table 8: Results of AlpacaEval 2 for Zephyr backbone. LC and WR denote length-controlled and raw win rate in percentage (%) respectively.
Method AlpacaEval 2
LC WR
Base 12.69 10.71
PPO 13.32 8.55
DPO 19.03 17.67
Appendix E Prompt used by LLM-as-a-judge
We provide the prompt used by LLM-as-a-judge in Figure 6.
LLM-as-a-Judge Prompt Template Review the user’s question and the corresponding response using the additive 5-point scoring system described below. Points are accumulated based on the satisfaction of each criterion: Add 1 point if the response is relevant and provides some information related to the user’s inquiry, even if it is incomplete or contains some irrelevant content. Add another point if the response addresses a substantial portion of the user’s question, but does not completely resolve the query or provide a direct answer. Award a third point if the response answers the basic elements of the user’s question in a useful way, regardless of whether it seems to have been written by an AI Assistant or if it has elements typically found in blogs or search results. Grant a fourth point if the response is clearly written from an AI Assistant’s perspective, addressing the user’s question directly and comprehensively, and is well-organized and helpful, even if there is slight room for improvement in clarity, conciseness or focus. Bestow a fifth point for a response that is impeccably tailored to the user’s question by an AI Assistant, without extraneous information, reflecting expert knowledge, and demonstrating a high-quality, engaging, and insightful answer. User: {instruction} Response: {response} After examining the user’s instruction and the response, provide brief step-by-step justifications and conclude with a score between 0 to 5. You must follow the format below: Evaluation: <<<evaluation>>> Score: <<<score>>>
Figure 6: We follow the prompt template used by Yuan et al. (2024) to use LLMs to judge model responses and construct paired dataset for further preference tuning.
Appendix F Hyperparameter tuning
In our experiments, the hyperparameters, including $\gamma$ and $\beta$, were selected based on the model performance on AlpacaEval 2 (AE2). These parameters were tuned for each method and model separately (with $\gamma$ specific to our approach) but remained fixed across iterations. Specifically, $\beta$ was tuned during the first iteration from the set {0.01, 0.1}, while $\gamma$ was tuned using the ranges {0.0, 0.25, 0.50, 0.75, 1.0} for the Zephyr backbone and {0.0, 0.10, 0.25, 0.50, 0.75, 1.0} for the Llama-3 backbone. The best hyperparameter setting identified in AlpacaEval 2 was directly applied in Arena-Hard (AH) for all methods and models without any further tuning.
Compared to baseline approaches, our approach has the additional hyperparameter $\gamma$ which may offer potential advantages during evaluation when hypertuning is performed on AE2. To address this concern, we treat the two benchmarks (AE2 and AH) as proxies for validation and test sets to ensure fair comparisons. Specifically, we validate on AE2 while testing on AH, and then reverse the roles of the two benchmarks. Below we show the results in Tables 9 and 10:
Table 9: Model performance on AlpacaEval 2 (AE2) and Arena-Hard (AH) across different $\gamma$ for the Zephyr Backbone. LC denotes the length-controlled win rate. CI denotes confidence interval.
$\gamma$ LC. AE2 LC.AH: Score (95% CI)
0.00 18.88 17.6 (-1.4, 1.7)
0.25 17.87 16.1 (-1.2, 1.4)
0.50 20.71 18.4 (-1.3, 1.6)
0.75 16.90 14.9 (-1.6, 1.4)
1.00 13.40 13.0 (-1.3, 1.3)
Table 10: Model performance on AlpacaEval 2 (AE2) and Arena-Hard (AH) across different $\gamma$ for the Llama-3 Backbone. LC denotes the length-controlled win rate. CI denotes confidence interval.
$\gamma$ LC. AE2 LC.AH: Score (95% CI)
0.00 26.93 39.1 (-2.3, 2.5)
0.10 27.55 39.1 (-2.0, 2.4)
0.25 24.78 33.3 (-2.1, 1.8)
0.50 23.59 32.2 (-2.3, 2.2)
0.75 22.57 27.4 (-2.0, 1.8)
1.00 22.50 24.5 (-1.7, 1.8)
The two benchmarks and two backbone models constitute four validation-test instances. We observe that the $\gamma$ selected from the validation set/benchmark consistently achieves the best performance on the test set across all instances, demonstrating the optimal value of $\gamma$ is robust to the choice of the validation set/benchmark. Meanwhile, we do observe that the optimal $\gamma$ for different backbone models is varied, which is expected, as different backbones produce responses of varying quality. Higher-quality generated data reduces the need for replayed experience, aligning with our observations for the Llama-3 backbone.